OpenStudy (anonymous):

NEED HELP ASAP!!! WILL GIVE MEDAL!! If the volume of a box is 180, what is the volume of the box if it is scaled down by a factor of 1/10?

OpenStudy (anonymous):

@selfconfidence777

OpenStudy (anonymous):

@coolman500

OpenStudy (anonymous):

@spongebob108

OpenStudy (anonymous):

@Jamierox4ev3r

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@rebeka

TheSmartOne (thesmartone):

Lets say this box is a cube.

OpenStudy (anonymous):

its a rectangular prism

TheSmartOne (thesmartone):

Well it doesn't say so, but to keep the question simpler, we can just make it a cube.

TheSmartOne (thesmartone):

So the volume of a cube = side x side x side

OpenStudy (anonymous):

yep

TheSmartOne (thesmartone):

So with what we are given: 180 = side^3

OpenStudy (anonymous):

@MrProphecyXII

TheSmartOne (thesmartone):

and they want to multiply it by a scale factor of 1/10

OpenStudy (anonymous):

they want me 2 scale it down

TheSmartOne (thesmartone):

That means: \(\sf\Large\frac{side}{10}\times\frac{side}{10}\times\frac{side}{10}=?\)

TheSmartOne (thesmartone):

But we already know that side x side x side is the volume which they gave us as 180

OpenStudy (anonymous):

yes

TheSmartOne (thesmartone):

So we have: \(\sf\Large\frac{side}{10}\times\frac{side}{10}\times\frac{side}{10}=\frac{side^3}{10\times10\times10}=\frac{180}{10^3}=?\)

OpenStudy (anonymous):

yes

TheSmartOne (thesmartone):

can you do the calculation?

OpenStudy (anonymous):

? ______ 1,000

TheSmartOne (thesmartone):

yup

OpenStudy (anonymous):

180 ______ 1,000

TheSmartOne (thesmartone):

and 180/1000 =?

OpenStudy (anonymous):

.0.18?

TheSmartOne (thesmartone):

and that is your final answer. :)

OpenStudy (anonymous):

yes it is!

OpenStudy (anonymous):

Thank you soooo much!!